OpenStudy (anonymous):

What is ( (tan^2)x)(2x)=3

OpenStudy (danjs):

\[2x \tan^{2}(x) = 3\] is that it?

OpenStudy (jhannybean):

Might be. I'm looking at it like O______________________o

OpenStudy (anonymous):

I think you have to use \(\cos ^2x = 1-\sin^2 x\)

OpenStudy (anonymous):

But even then I'm not completely sure of the trick here.

Miracrown (miracrown):

Are we trying to solve for x in this problem?

OpenStudy (mathmate):

It's a non-linear problem, and is not likely to have an elementary solution. Try numerical methods.

OpenStudy (dumbcow):

@Asiah321 could you confirm the equation? what @DanJS posted can not be solved algebraically

Miracrown (miracrown):

\[(\tan ^{2}x) \space (2x) \space = 3\]

Miracrown (miracrown):

\[(tanx)^{2} \space (2x) \space =3\]

Miracrown (miracrown):

You can't actually solve this by hand because you can't get x alone tan^2 x means (tanx)^2 If you want to get x, you need a calculator You would either have to graph y = (tanx)^2*2x and y = 3 and find where they intersect Or graph y = (tanx)^2*2x - 3 and find where it hits the x-axis

OpenStudy (mathmate):

There are 4 roots between 0 & 2pi. The first one is around 0.9093 rad.

OpenStudy (mathmate):

Newton's method would come handy for this, or even bisection method.

Miracrown (miracrown):

Right

Miracrown (miracrown):

I agree with that too. Whichever methods allow you to approximate roots

Miracrown (miracrown):

Newton's method might be a drag though since you would need to do it for each root

OpenStudy (danjs):

i just guessed that was his question, i could be wrong

Miracrown (miracrown):

Just using a calculator straight out seems to be the easiest method here You would need a calculator for newton's method in any case

OpenStudy (mathmate):

For those interested, I graphed it as suggeste by @Miracrown, it gives roots at 0.9,2.5,3.7,and 5.8 between 0 and 2pi, for those who would give it a try.

Miracrown (miracrown):

Yes, I agree ^

OpenStudy (mathmate):

and for the Newton fans, y'(x)=2tan^2(x)+4x tan(x)/cos(x)^2

OpenStudy (mathmate):

@DanJS we're all in the dark, and you're the only one with a candle. :) It's up to @Asiah321 to correct your best interpretation if it's not correct.